True or False. Explain your answer.
The ordered pair(1,2)is the solution set to the equation
2x+y=4

Found 2 solutions by tutorcecilia, AnlytcPhil:
Answer by tutorcecilia(2152)

2(1)+(2)=4 [Plug-in the given values for (x, y)into the equation]
2+2=4 [Solve]
4=4 [The answer is true. When the values for (x, y) are plugged into the equation, it causes both sides to become equal]

Answer by AnlytcPhil(1810)

True or False. Explain your answer. 
The ordered pair(1,2)is the solution set to the equation
2x+y=4


Let's see if it is a solution

the first number in the ordered pair is the value to
substitute for x.

the second number in the ordered pair is the value to
substitute for y.

So we substitute 1 for x and 2 for y in 

       2x + y = 4

to see if we get two equal numbers on the two sides
    
     2(1) + 2 = 4
        2 + 2 = 4
            4 = 4

Yes this is A solution to the equation 
 
       2x + y = 4

What bothers me is your words "THE SOLUTION" instead
of "A SOLUTION".  There are many solutions to this
equation.  For instance 

(-2,8) is another solution because

       2x + y = 4
      
    2(-2) + 8 = 4
       -4 + 8 = 4
            4 = 4

(3, -2) is yet another solution because

       2x + y = 4
      
  2(3) + (-2) = 4
        6 - 2 = 4
            4 = 4

(2, 0) is yet another solution because

       2x + y = 4
      
   2(2) + (0) = 4
        4 - 0 = 4
            4 = 4

(4, -4) is yet another solution because

       2x + y = 4
      
  2(4) + (-4) = 4
        8 - 4 = 4
            4 = 4

(-4, 12) is yet another solution because

       2x + y = 4
      
2(-4) + (-12) = 4
       -6 - 2 = 4
            4 = 4

So, you see, there are many many solutions
to the equation

       2x + y = 4
      
not just one.

Edwin